# Hybrid Quantum Physics-informed Neural Network: Towards Efficient Learning of High-speed Flows

### v1

@@ Abstract
This study assesses the potential use of hybrid quantum physics-informed neural network (HQPINN) to model high-speed flows as an alternative to classical PINN and quantum neural network options. The model integrates parameterized quantum circuit (PQC) with classical neural network in parallel as input to a physics-based optimization. For problems with harmonic solutions, the HQPINN exhibits superior accuracy and trainability compared to both classical and quantum models at low parameter costs. For transonic flows, the hybrid approach yields modest results and additionally suffers from poor trainability if the quantum layer were under-parameterized. Our results highlight inherent limitations in deploying quantum neural networks in PINN applications, and potential use of hybrid architectures as a general tool in problems where the nature of the solution is not known a-priori.
@@ 1 Introduction
Prediction of compressible air flow features at transonic speed is critical to efficient aerofoil design. This problem involves solving a nonlinear set of partial differential equations (PDEs) in a domain bound by an irregular aerofoil geometry, aspects which pose formidable challenge to traditional computational fluid dynamics (CFD) [Sandeep2021]. In recent years, machine learning techniques, particularly physics-informed neural networks (PINNs) [Raissi2019], have emerged as a promising tool for handling high-speed flows, with certain advantages in data-driven neural networks and inverse applications [Mao2020, Jagtap2022, Wassing2024, Cao2024].
@@
Conceptually, PINNs use differentiable neural networks to approximate the solution of some governing equations, in a way that obeys imposed physical laws, including nonlinear phenomena and geometric constraints [Raissi2019]. This approach allows for the incorporation of prior physical knowledge directly into the learning process, enhancing the model’s ability to generalize from limited data. However, as a neural network strategy, PINNs are also limited by extensive use of automatic differentiation for training, which often leads to long training times and prohibitive computational costs [Cao2024].
@@
Quantum computing [nielsen2010quantum] offers a promising avenue to accelerate scientific computing and machine learning, particularly through quantum machine learning (QML) [Setty_2025]. QML performs machine learning tasks on a parameterized quantum circuit (PQC), also known as quantum neural network, which exploits quantum features such as superposition and entanglement for higher trainability and expressivity [Abbas2021]. PQCs are not only universal approximators like classical multi-layered perceptrons (MLPs), but they also act as truncated Fourier series models [schuld2021effect]. This unique feature enables PQCs to capture complex patterns and relationships in data more efficiently than classical models.
@@
Recent studies have seen efforts to replace classical neural networks with quantum PQC, either partially or entirely. For instance, Arthur and Date [arthur2022] proposed a hybrid quantum neural network consisting of a quantum PQC layer and a classical MLP layer in series for classification problems showing a higher accuracy compared to pure PQC. Kordzanganeh et al. [Kordzanganeh2023exponential] proposed a parallel hybrid quantum neural networkof extracting both harmonic and non-harmonic features from a dataset [Kordzanganeh2023]. Since then, there are studies of hybrid classical-quantum neural networks modeling flows past a cylinder [Jaderberg2023] and mixing flows via transfer learning [Sedykh_2024]. More recently, Xiao et al. [xiao2024physics] simulated quantum neural networks for partial differential equations, which can deliver high accuracy solutions using fewer trainable parameters.
@@
In this study, we explore the use of hybrid quantum PINN (HQPINN) on high-speed flows problems. This is a challenging CFD problem where pure quantum neural networks are not favored due to non-harmonic solutions that include shocks and discontinuities [xiao2024physics]. We argue that, in the absence of detailed prior knowledge of the problem, a hybrid quantum PINN could flexibly model flows that include both smooth and discontinuous features. For verification, we assess the use of HQPINNs on model problems by benchmarking their trainability and training accuracy against classical PINNs and PQCs with varying complexities.
@@ 2.1 Governing equations
For 2D inviscid compressible flow, the governing Euler equations are [Mao2020, Jagtap2022]:
@@
$\rho$ is the fluid density, $(u,v)$ are velocity components in $(x,y)$, $E$ is the total energy and $p$ is the pressure. The equation of state for a polytropic gas is
@@ 2.3 Parameterized quantum circuit
Quantum machine learning (QML) is the application of quantum computing to machine learning tasks [Benedetti2019]. In the literature, QML circuits are also often referred to as variational quantum circuits, quantum circuit learning, quantum neural networks or parameterized quantum circuits (PQC).
@@
where $|\psi(\boldsymbol{\theta},\boldsymbol{\varphi})\rangle$ is the quantum state at the point of measurement. Note that the quantum model fits a truncated Fourier series function [schuld2021effect], as
@@
Like a multi-layered perceptron (MLP), a PQC is also an universal approximator that can functionally replace at least parts of a neural network, forming a classical-quantum interface known as hybrid quantum neural network (HQNN) [Kordzanganeh2023]. There are a few reasons why this might be a good idea. Quantum computers already offer exponential computational capacity over classical computers in the form of a large Hilbert space and may also enhance the machine learning process [Melnikov2023]. Also, the tools of differential quantum circuits [Kyriienko2021] align with classical neural networks using automatic differentiation and back-propagation. Finally, the Fourier model of a PQC (Eq. 9) could extract harmonic features from a suitable dataset and complement a classical neutral network [Kordzanganeh2023].
@@ 2.4 Hybrid quantum physics-informed neural networks
Here, we propose a hybrid quantum physics-informed neural network (HQPINN) as sketched in Figure 1. The HQPINN samples $(x,t)$ data and output $\{\rho,u,p\}$ using a hybrid architecture with quantum and classical layers arranged in parallel [Kordzanganeh2023]:
Figure 1: Schematic of proposed HQPINN which samples $(x,t)$ data into a parallel architecture coupling a quantum layer (top), an $n$ qubit PQC formed by alternating $l-1$ layers of feature map $S(\boldsymbol{\varphi})$ and $l$ layers of parameterized ansatz $A(\boldsymbol{\theta})$, with a classical layer (bottom), a neural network formed by an input layer with $n$ nodes and $L$ hidden layers of $N$ nodes [Kordzanganeh2023]. The quantum output $q_{[1,n]}$ and classical output $c_{[1,n]}$ combine linearly to approximate $(\rho,u,p)$ and their informed functions $(f(\cdot),\partial_{t},\nabla\cdot)$, both collectively contribute towards a loss function for optimization [Mao2020].
@@
The final neural network layer is a weighted linear combination of the outputs from the quantum and classical layers $o_{i}=s_{i}^{c}c_{i}+s_{i}^{q}q_{i}$ [Kordzanganeh2023exponential] to produce an output vector $\{\rho,u,p\}$ for minimization of the loss function Eq. 5. We tested the hybrid classical–quantum neural network (Fig. 1) against a classical–classical neural network where the parallel network layers are both classical layers, and a quantum–quantum neural network where the parallel network layers are both quantum layers. For the introductory 1D harmonic oscillator problem, the HQPINN was found to outperform both classical–classical and quantum–quantum neural network architectures (Appendix A.2). Henceforth, we aim to verify if the hybrid classical-quantum scheme holds up for more challenging problems such as high-speed shock flows.
@@ 3 Numerical results
We test and compare classical and quantum PINN models based on two example cases on high-speed flows, one with smooth solutions ([Mao2020], example 6) and the other with discontinuous solutions ([Mao2020], example 1). Here we consider three models:
@@
For convenience, we adopt the following itemized notation for each model. The label cc denotes classical model, followed by the number of nodes per hidden layer and the number of hidden layers respectively. The label qq denotes quantum model, followed by number of trainable layers. The label hy denotes hybrid classical–quantum model, followed by the number of nodes per hidden layer, the number of hidden layers and number of trainable quantum layers respectively.
Models are implemented on Pytorch [Paszke2019] interfaced with Pennylane [Bergholm2018] and run on distributed CPU without GPU acceleration. Loss functions (Eq. 5) are minimized via mean square error (Eq. 6 and 7) using Adam optimizer for 20 000 steps with learning rate 0.0005. The quantum layer is simulated classically, which results in an actual compute time that is slower than a classical layer by at least an order of magnitude[xiao2024physics].
3.1 Smooth Euler
Consider a 1D Euler equation with smooth solution ([Mao2020], example 6). We use periodic boundary conditions and initial condition,
@@ 3.1 Smooth Euler
in a domain defined by $x\in(-1,1)$ and $t\in(0,2)$. Following ([Mao2020], example 6), we randomly sample $N_{ic}=50$, $N_{bc}=50$ and $N_{F}=2000$ training points for initial condition $ic$, boundary condition $bc$ (Eq. 6) and domain $F$ (Eq. 7) respectively.
Figure 2 shows a traveling wave solution of the density based on classical cc-10-4 model, with high fidelity agreement with the exact solution indicated by their normalized deviation. Table 1 compares the number of trainable parameters required for each model and the resultant loss and error at the 20 000th epoch. Here, the best model is the quantum qq-2 model with an average loss of $2.63\times 10^{-8}$ using only 87 classical and 36 quantum trainable parameters, substantially lower than classical models with comparable performance. Larger quantum models such as qq-4 underperform due to poor trainability arising from excessive expressibility for a simple solution, a phenomenon also observed in QML elsewhere [xiao2024physics]. Interestingly, the best hybrid model hy-10-4-2 is not a combination of the best classical cc-10-7 and quantum qq-2 models.
@@
Model Size Trainable parameters Loss Density error Pressure error
cc 10–4 845 $6.13\times 10^{-7}$ $1.59\times 10^{-6}$ $3.28\times 10^{-9}$
10–7 1505 $\mathbf{3.16\times 10^{-7}}$ $\mathbf{7.33\times 10^{-7}}$ $\mathbf{2.22\times 10^{-9}}$
20–4 2845 $6.65\times 10^{-7}$ $2.02\times 10^{-6}$ $3.96\times 10^{-9}$
qq 2 87(+36) $\mathbf{2.63\times 10^{-8}}$ $\mathbf{9.76\times 10^{-9}}$ $\mathbf{1.01\times 10^{-9}}$
3 87(+54) $1.71\times 10^{-7}$ $1.34\times 10^{-7}$ $1.87\times 10^{-9}$
4 87(+72) $2.75\times 10^{-7}$ $2.80\times 10^{-7}$ $2.46\times 10^{-9}$
hy 10–4–2 448(+36) $\mathbf{1.98\times 10^{-7}}$ $\mathbf{2.29\times 10^{-7}}$ $8.44\times 10^{-9}$
10–7–2 778(+36) $3.31\times 10^{-7}$ $3.24\times 10^{-7}$ $9.04\times 10^{-9}$
20–4–2 1448(+36) $8.62\times 10^{-7}$ $3.19\times 10^{-7}$ $\mathbf{4.98\times 10^{-9}}$
Table 1: Trainable parameter count for each PINN model with additional quantum parameters in parenthesis for smooth solution problem (Section 3.1). Mean loss and relative error for density and pressure at the 20 000th epoch, averaged over 6 runs. Best results shown in bold for each model.
Figure 3 shows training loss against epochs for selected neural network models with shallow, deep or wide characteristics. Our results show that classical models train fast but with low accuracy, in contrast to quantum models which train slower but with high accuracy. Hybrid models inherit high trainability from the classical layer and high accuracy from quantum layer, the best features from each model. The exception is the deep hy-10-7-2 model (Fig. 3c), where training is weighted towards the classical layer than the quantum layer.
@@
3.2 Discontinuous Euler
Consider a 1D Euler equation with moving contact discontinuity ([Mao2020], example 1). We use Dirichlet boundary conditions,
@@ 3.2 Discontinuous Euler
in a domain defined by $x\in(0,1)$ and $t\in(0,2)$. Following ([Mao2020], example 1), we randomly sample $N_{ic}=60$, $N_{bc}=60$ and $N_{F}=1000$ training points for initial condition $ic$, boundary condition $bc$ and domain $F$ respectively.
Figure 4 shows moving contact discontinuity in the density based on classical cc-10-4 model, where the normalized deviation from exact solution concentrated along the discontinuity itself. Table 2 compares the trainable parameters required for each model and the resultant loss and error at the 20 000th epoch. Here, the best model is the shallow classical cc-10-4 model with an average loss of $6.89\times 10^{-7}$ for 845 trainable parameters. Due to the discontinuous non-harmonic solution, the quantum qq models significantly underperformed compared to classical cc models, but improves with increasing number of trainable parameters. Overall, we find the hybrid models performed slightly worse than classical models, but significantly better than quantum models.
@@
Model Size Trainable parameters Loss Density error Pressure error
cc 10–4 845 $\mathbf{6.89\times 10^{-7}}$ $\mathbf{2.89\times 10^{-4}}$ $1.92\times 10^{-6}$
10–7 1505 $1.39\times 10^{-6}$ $4.07\times 10^{-4}$ $\mathbf{6.06\times 10^{-7}}$
20–4 2845 $1.51\times 10^{-6}$ $4.67\times 10^{-4}$ $6.59\times 10^{-7}$
qq 2 87(+36) $5.00\times 10^{-4}$ $1.20\times 10^{-3}$ $4.51\times 10^{-5}$
3 87(+54) $1.45\times 10^{-4}$ $1.30\times 10^{-3}$ $1.53\times 10^{-5}$
4 87(+72) $\mathbf{1.39\times 10^{-4}}$ $\mathbf{9.65\times 10^{-4}}$ $\mathbf{1.24\times 10^{-5}}$
hy 10–4–2 448(+36) $7.53\times 10^{-6}$ $5.29\times 10^{-4}$ $\mathbf{1.96\times 10^{-6}}$
10–7–2 778(+36) $3.93\times 10^{-5}$ $3.55\times 10^{-4}$ $3.52\times 10^{-5}$
20–4–2 1448(+36) $\mathbf{6.98\times 10^{-6}}$ $\mathbf{2.37\times 10^{-4}}$ $3.29\times 10^{-6}$
Table 2: Trainable parameter count for each PINN model with additional quantum parameters in parenthesis for discontinuous problem (Section 3.2). Mean loss and relative error for density and pressure averaged over 6 runs at the 20 000th epoch. Best results in bold for each model.
Figure 5 shows density profiles at time $t=2$ where we see the quantum model fails to capture the moving contact discontinuity with an oscillatory waveform with overshoot. The hybrid model is able to correct for the perturbation error with the parallel classical layer and achieve a solution fidelity closer to that of the classical model.
@@
Figure 6 shows training loss against epochs for selected neural network models with shallow, deep or wide characteristics. The classical models outperformed the quantum models in all cases, while the hybrid models suffer from significant variance in both trainability and accuracy in results. We find that hybrid hy-10-4-2 and hy-10-7-2 models are unable to match the respective classical models, especially the deep cc-10-7 model in terms of trainability. However, the wide hybrid hy-20-4-2 model is able to match the wide classical cc-20-4 model, the latter likely over-parameterized.
@@
We test HQPINN on a 2D steady Euler equation (Eq. 3) for transonic flow past an NACA0012 aerofoil with chord $(0,1)$ set on a square Cartesian grid $x\in(-1,3.5),y\in(-2.25,2.25)$. Setting $U=(\rho,u,v,T)$ for density, horizontal and vertical velocities, and temperature, we apply an inlet boundary condition $U_{in}=(\rho_{in},u_{in},v_{in},T_{in})=(1.225,272.15,0.0,288.15)$, outlet boundary condition $P_{out}=0$ and periodic boundary conditions on the side boundaries, all in SI units.
@@
It should be pointed out that the transonic shock problem is, in fact, highly challenging even for vanilla PINNs. It is only until recently that add-on strategies, such as coordinate transformation, artificial viscosity, adaptive weights and shock sensors [wassing2024physics, Ren2024, McClenny2023], enable PINNs to handle shock regions. While exploring these strategies are beyond the scope of this work, to mitigate the expression of shocks, we apply an adaptive gradient-weight [Liu2024],
@@
The transonic flow requires a minimum 40 nodes per hidden layer to resolve (Table 3). In this regard, the quantum models considered are severely under-parameterized. Yet, it is currently challenging to study the potential of larger quantum models as the time required to simulate them classically scales unfavorably, as shown by the wall time. While this bottleneck may be relieved in the future by the development of fault-tolerant quantum hardware and software acceleration, we take this opportunity to study the effects of unbalanced parameterization between the classical and quantum layers of the hybrid model.
Model Size Trainable parameters Time (s/epoch) Loss Pressure error
cc 40–4 10 628 0.11 $1.11\times 10^{-2}$ $1.35\times 10^{-3}$
40–7 20 468 0.21 $2.48\times 10^{-3}$ $6.79\times 10^{-4}$
80–4 40 388 0.20 $3.13\times 10^{-3}$ $7.74\times 10^{-4}$
qq 2 140(+48) 1.9 $9.67\times 10^{1}$ $3.59\times 10^{-3}$
4 140(+96) 4.6 $8.48\times 10^{1}$ $2.53\times 10^{-3}$
6 140(+144) 9.0 $6.65\times 10^{1}$ $4.21\times 10^{-3}$
hy 40–4–2 5360(+48) 1.3 $1.36\times 10^{-1}$ $4.05\times 10^{-3}$
40–7–2 10 280(+48) 1.1 $3.71\times 10^{0}$ $1.95\times 10^{-3}$
80–4–2 20 240(+48) 1.1 $2.56\times 10^{-2}$ $2.46\times 10^{-3}$
Table 3: Trainable parameter count for each PINN model with additional quantum parameters in parenthesis for discontinuous problem (Section 3.3). Average wall time shown in seconds per epoch. Mean loss and relative error for pressure averaged over 6 runs at the 42 000th epoch.
Figure 8a show that larger quantum models do not scale well in trainability and accuracy. This may change with further increase in the number of trainable parameters. Hybrid models underperform compared to classical models (Fig. 8b-d). Unlike in the previous example (Fig. 7), hybrid models for the transonic flow problem display significantly poorer trainability compared to their classical counterparts, due to the under-parameterized quantum layer.

### v2

@@ Abstract
This study benchmarks hybrid quantum physics-informed neural network (HQPINN) to model high-speed flows, compared against classical physics-informed neural networks (PINNs) and fully quantum neural networks (QNNs). The HQPINN architecture integrates a parameterized quantum circuit (PQC) with a classical neural network in parallel, trained via a physics-informed loss. Across harmonic, non-harmonic, and transonic benchmarks, HQPINNs demonstrate balanced performance, offering competitive accuracy and stability with reduced parameter cost. Quantum PINNs are highly efficient for harmonic problems achieving the lowest loss with minimal parameters due to their Fourier structure, but struggle to generalize in non-harmonic settings involving shocks and discontinuities. HQPINNs mitigate such artifacts, and with sufficient parameterization, can match the performance of classical models in more complex regimes. Although constrained by current quantum emulation costs and scalability, HQPINNs show promise as general-purpose solvers, offering parameter efficiency with robust fallback behavior, particularly suited for problems where the nature of the solution is not known a-priori.
Keywords: Hybrid Quantum-Classical Computing; Physics-Informed Neural Networks; High-speed Flows; Shock Capturing; Quantum Neural Networks
@@ 1 Introduction
Prediction of compressible air flow features at transonic speed is critical to efficient aerofoil design. This problem involves solving a nonlinear set of partial differential equations (PDEs) in a domain bound by an irregular aerofoil geometry, aspects which pose formidable challenge to traditional computational fluid dynamics (CFD) [1]. In recent years, machine learning techniques, particularly physics-informed neural networks (PINNs) [2], have emerged as a promising tool for handling high-speed flows, with certain advantages in data-driven neural networks and inverse applications [3, 4, 5, 6].
@@
Conceptually, PINNs use differentiable neural networks to approximate the solution of some governing equations, in a way that obeys imposed physical laws, including nonlinear phenomena and geometric constraints [2]. This approach allows for the incorporation of prior physical knowledge directly into the learning process, enhancing the model’s ability to generalize from limited data. However, as a neural network strategy, PINNs are also limited by extensive use of automatic differentiation for training, which often leads to long training times and prohibitive computational costs [6].
@@
Quantum computing [7] offers a promising avenue to accelerate scientific computing and machine learning, particularly through quantum machine learning (QML) [8]. QML performs machine learning tasks on a parameterized quantum circuit (PQC), also known as quantum neural network, which exploits quantum features such as superposition and entanglement for higher trainability and expressivity [9]. PQCs are not only universal approximators like classical multi-layered perceptrons (MLPs), but they also act as truncated Fourier series models [10]. This unique feature enables PQCs to capture complex patterns and relationships in data more efficiently than classical models.
@@
Recent studies have seen efforts to replace classical neural networks with quantum PQC, either partially or entirely [11, 12, 13, 14, 14, 15, 16, 17, 18, 19, 20]. For instance, Arthur and Date [11] proposed a hybrid quantum neural network consisting of a quantum PQC layer and a classical MLP layer in series for classification problems showing a higher accuracy compared to pure PQC. Kordzanganeh et al. [12] proposed a parallel hybrid quantum neural network capable of extracting both harmonic and non-harmonic features from a dataset [13]. Since then, there are studies of hybrid classical-quantum neural networks modeling flows past a cylinder [14] and mixing flows via transfer learning [15]. More recently, Xiao et al. [16] and Siegl et al. [17] simulated quantum neural networks for PDEs, which can deliver high accuracy solutions using fewer trainable parameters [18, 19]. Berger et al. [20] leveraged on feedforward neural networks as embedding functions to solve nonlinear PDEs. These studies suggest that hybrid quantum-classical neural network architectures may offer advantages such as enhanced expressivity, in handling more complex PDEs but thus far, work in this area has been limited to problems with smooth solutions [16].
@@
In this study, by leveraging the benefits of hybrid quantum-classical neural network architectures, we conduct a systematic benchmarking study of HQPINNs on problems ranging from smooth solutions to discontinuous flows and transonic shocks. This is a challenging CFD problem where pure quantum neural networks are not favored due to non-harmonic solutions that include shocks and discontinuities [16, 17]. To the best of our knowledge, this is the first time HQPINN is applied to high speed flow problems to assess potential of classical and quantum PINNs. We evaluate when hybrid architectures provide benefit, when they fail, and what architectural or training limitations contribute to those outcomes. For verification, we assess the use of HQPINNs on model problems by benchmarking their trainability and training accuracy against classical PINNs and PQCs with varying complexities.
@@ 2.1 Governing equations
For 2D inviscid compressible flow, the governing Euler equations are [3, 4]:
@@
[…]
@@ 2.3 Parameterized quantum circuit
Quantum machine learning (QML) is the application of quantum computing to machine learning tasks [21]. In the literature, QML circuits are also often referred to as variational quantum circuits, quantum circuit learning, quantum neural networks or parameterized quantum circuits (PQC).
@@
where $|\psi(\boldsymbol{\theta},\boldsymbol{\varphi})\rangle$ is the quantum state at the point of measurement. Note that the quantum model fits a truncated Fourier series function [10], as
@@
Like a multi-layered perceptron (MLP), a PQC is also an universal approximator that can functionally replace at least parts of a neural network, forming a classical-quantum interface known as hybrid quantum neural network (HQNN) [13]. There are a few reasons why this might be a good idea. Quantum computers already offer exponential computational capacity over classical computers in the form of a large Hilbert space and may also enhance the machine learning process [22]. Also, the tools of differential quantum circuits [23] align with classical neural networks using automatic differentiation and back-propagation. Finally, the Fourier model of a PQC (Eq. 9) could extract harmonic features from a suitable dataset and complement a classical neutral network [13].
@@ 2.4 Hybrid quantum physics-informed neural networks
Here, we propose a hybrid quantum physics-informed neural network (HQPINN) as sketched in Figure 1. The HQPINN samples $(x,t)$ data and output $\{\rho,u,p\}$ using a hybrid architecture with quantum and classical layers arranged in parallel [13]:
Figure 1: Schematic of proposed HQPINN which samples $(x,t)$ data into a parallel architecture coupling a quantum layer (top), an $n$ qubit PQC formed by alternating $l-1$ layers of feature map $S(\boldsymbol{\varphi})$ and $l$ layers of parameterized ansatz $A(\boldsymbol{\theta})$, with a classical layer (bottom), a neural network formed by an input layer with $n$ nodes and $L$ hidden layers of $N$ nodes [13]. The quantum output $q_{[1,n]}$ and classical output $c_{[1,n]}$ combine linearly to approximate $(\rho,u,p)$ and their informed functions $(f(\cdot),\partial_{t},\nabla\cdot)$, both collectively contribute towards a loss function for optimization [3].
@@
The final neural network layer is a weighted linear combination of the outputs from the quantum and classical layers $o_{i}=s_{i}^{c}c_{i}+s_{i}^{q}q_{i}$ [12], where $s_{i}^{c}$ and $s_{i}^{q}$ are scalar weights applied to each output component. In this study, we fix $s_{i}^{c}=s_{i}^{q}=1$, but future work may consider learnable fusion layers or nonlinear mappings such as $o_{i}=\mathrm{MLP}([c_{i},q_{i}])$. The combined output is interpreted as predicted physical state $\{\rho,u,p\}$ for minimization of the loss function Eq. 5.
We tested the hybrid classical–quantum neural network (Fig. 1) against a classical–classical neural network where the parallel network layers are both classical layers, and a quantum–quantum neural network where the parallel network layers are both quantum layers. For the introductory 1D harmonic oscillator problem, the HQPINN was found to outperform both classical–classical and quantum–quantum neural network architectures (Appendix A.2). Henceforth, we aim to verify if the hybrid classical-quantum scheme holds up for more challenging problems such as high-speed shock flows.
@@ 3 Numerical results
We benchmark classical and quantum PINN models based on two example cases on high-speed flows, one with smooth solutions ([3], example 6) and the other with discontinuous solutions ([3], example 1). Here we consider three models:
@@
To facilitate direct comparison between models, we adopt the following structured notion:
cc-N-L
: Classical network with L layers of N nodes.
qq-l
: Quantum network with l PQC layers with number of qubits fixed by number of outputs.
hy-N-L-l
Hybrid network with classical layer of size NL and quantum layer with $l$ PQC layers.
Models are implemented on Pytorch [24] interfaced with Pennylane [25] and run on distributed CPU without GPU acceleration. Loss functions (Eq. 5) are minimized via mean square error (Eq. 6 and 7) using Adam optimizer for 20 000 steps with learning rate 0.0005. Quantum layers are simulated classically, which imposes a significant time overhead compared to classical layers, by at least two orders of magnitude for the same number of trainable parameters [16, 17].
3.1 Smooth Euler Equation (Harmonic Regime)
Consider a 1D Euler equation with smooth solution ([3], example 6). We use periodic boundary conditions and initial condition,
@@ 3.1 Smooth Euler Equation (Harmonic Regime)
in a domain defined by $x\in(-1,1)$ and $t\in(0,2)$. Following ([3], example 6), we randomly sample $N_{ic}=50$, $N_{bc}=50$ and $N_{F}=2000$ training points for initial condition $ic$, boundary condition $bc$ (Eq. 6) and domain $F$ (Eq. 7) respectively.
Figure 2 shows a traveling wave solution of the density based on classical cc-10-4 model, with high fidelity agreement with the exact solution indicated by their normalized deviation. Table A1 (Appendix A.5) compares the number of trainable parameters required for each model and the resultant loss and error at the 20 000th epoch.
Here, the best model is the quantum qq-2 model with an average loss of $2.63\times 10^{-8}$, a tenfold improvement compared to the best classical cc-10-7 model with an average loss of $3.16\times 10^{-7}$. In addition, the quantum qq-2 model requires only 87 classical and 36 quantum trainable parameters vastly outperforming the best classical cc-10-7 model. However, note that larger quantum models suffer from poor trainability due to excessive expressibility, a phenomenon also observed in QML elsewhere [16].
The hybrid model hy-10-4-2 achieves strong balanced performance with high accuracy and trainability, as well as reduced parameter count (484 vs. 1505). Interestingly, the best hybrid model hy-10-4-2 is not a combination of the best classical cc-10-7 and quantum qq-2 models.
@@
Figure 3 shows training loss against epochs for selected neural network models with shallow, deep or wide characteristics. Our results show that classical models train fast but with low accuracy, in contrast to quantum models which train slower but with high accuracy. Hybrid models inherit high trainability from the classical layer and superior accuracy from quantum layer, the best features from each model. The exception is the deep hy-10-7-2 model (Fig. 3c), where training is weighted towards the classical layer than the quantum layer.
@@
3.2 Discontinuous Euler Equation (Contact Discontinuity)
Consider a 1D Euler equation with moving contact discontinuity ([3], example 1). We use Dirichlet boundary conditions,
@@ 3.2 Discontinuous Euler Equation (Contact Discontinuity)
in a domain defined by $x\in(0,1)$ and $t\in(0,2)$. Following ([3], example 1), we randomly sample $N_{ic}=60$, $N_{bc}=60$ and $N_{F}=1000$ training points for initial condition $ic$, boundary condition $bc$ and domain $F$ respectively.
Figure 4 shows moving contact discontinuity in the density based on classical cc-10-4 model, where the normalized deviation from exact solution concentrated along the discontinuity itself. Table A2 (Appendix A.5) compares the trainable parameters required for each model and the resultant loss and error at the 20 000th epoch. The shallow classical cc-10-4 model is the best model with an average loss of $6.89\times 10^{-7}$ for 845 trainable parameters. In contrast, quantum qq models significantly underperformed with average loss ¯$\overline{\mathcal{L}}_{qq}\sim\mathcal{O}(10^{-4})$, but which improves with increasing number of trainable quantum parameters (average loss decreases by 72% from qq-2 to qq-4). Overall, the hybrid models (¯$\overline{\mathcal{L}}_{hy}\sim\mathcal{O}(10^{-6}-10^{-5})$) performed slightly worse than classical models (¯$\overline{\mathcal{L}}_{cc}\sim\mathcal{O}(10^{-7}-10^{-6})$), but significantly better than quantum models (¯$\overline{\mathcal{L}}_{qq}\sim\mathcal{O}(10^{-4})$).
@@
Figure 5 shows density profiles at time $t=2$ where we see the quantum model exhibits spurious oscillations and fails to capture the shock discontinuity. In contrast, hybrid models are able to reduce overshoot observed in quantum models and achieve high solution fidelities similar to classical models.
@@
Figure 6 shows training loss against epochs for selected neural network models with shallow, deep or wide characteristics. We found that the classical models outperformed the quantum models in all cases, while the hybrid models suffer from significant variance in both accuracy (loss $\mathcal{L}$) and trainability (rate of change of $\mathcal{L}$). Figure 6a-c shows that hybrid models are unable to match the respective classical models in terms of loss accuracy. In addition, the deep hybrid hy-10-7-2 model performs worse than deep classical cc-hy-10-7 in terms of trainability (Fig. 6c). Importantly, note that the wide hybrid hy-20-4-2 model performs as well as the wide classical cc-20-4 model (Fig. 6d), while requiring only about 55% of the number of trainable parameters. This suggests that hybrid models may be advantageous for problems with non-harmonic solution under over-parameterized conditions.
@@
We test HQPINN on a 2D steady Euler equation (Eq. 3) for transonic flow past an NACA0012 aerofoil with chord (0,1)(0,1)( 0 , 1 ) set on a square Cartesian grid $x\in(-1,3.5),y\in(-2.25,2.25)$. Setting $U=(\rho,u,v,T)$ for density, horizontal and vertical velocities, and temperature, we apply an inlet boundary condition $U_{in}=(\rho_{in},u_{in},v_{in},T_{in})=(1.225,272.15,0.0,288.15)$, outlet boundary condition $P_{out}=0$ and periodic boundary conditions on the side boundaries, all in SI units.
@@
It should be pointed out that the transonic shock problem is, in fact, highly challenging even for vanilla PINNs. It is only until recently that add-on strategies, such as coordinate transformation, artificial viscosity, adaptive weights and shock sensors [26, 27, 28], enable PINNs to handle shock regions. While exploring these strategies are beyond the scope of this work, to mitigate the expression of shocks, we apply an adaptive gradient-weight [29],
@@
Because the transonic flow requires a minimum of 40 nodes per hidden layer to resolve, quantum models are significantly under-parameterized in comparison (Table A3, Appendix A.5). Figure 8 shows that under-parameterized quantum models do not scale in trainability and accuracy, which in turn also limits the trainability of hybrid models.
It is currently challenging to study the performance of larger quantum models as the time required to simulate them classically scales unfavorably. This bottleneck may be relieved in the future as fault-tolerant quantum hardware or accelerated hybrid software scheme become available.